OpenStudy (anonymous):

A trapezoid has an area of 4 square units and a height of 1 unit what are the possible whole number lengths for the bases

OpenStudy (anonymous):

Use the formula for the area of a trapezoid (A) with bases (b1 and b2) and height (h). A=(1/2)(b1+b2)(h) We know: A=4, h=1. Plug in those values. 4=1/2(b1+b2)(1) 4=(b1+b2)/2 8=b1+b2 b2=8-b1 Now, we can have these possible values for b1 and b2. b1=1, b2=8-1=7 b1=2, b2=8-2=6 b1=3, b2=8-3=5 b1=4, b2=8-4=4<---This is eliminated because the bases cannot be equal. If they were equal, we would have a rectangle, not a trapezoid. The rest are just the above with b1 and b2 switched. So it is unnecessary to go on. Thus, in units, the possible whole number lengths for the bases are: (1,7) (2,6) (3,5)